what is the mass of photon ? According to de broglie equation it is almost zero ….... but according to relativity, the mass of a particle moving at the speed of light is infinity(consider the particle as photon here ) . What is the real mass of electron ? pls explain how it is derived .

Grade:12

3 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The photon is massless. It has energy but no mass.
The de Broglie equation for a relativisitic particleis given by:
\lambda=\frac{h}{mv}=\frac{h}{\sqrt{2mK.E.}}=\frac{hc}{\sqrt{2m_{0}c^{2}K.E.}}
The rest mass is zero for the photon as the velocity of the particle approaches the velocity of light.
The real mass of electron is approximately9.1\times 10^{-31} kg
It is derived from the charge by mass ratio experiment.
Regards
Sumit
Aziz Alam IIT Roorkee
9 years ago
Mass, according to theory of relativity , depends upon the speed with which it is moving with respect to a particular frame of reference.Theory of relativity is given in accordaqnce with Maxwell’s theory of Electrodynamics and an assupmtion that mass of photon when it is at rest is zero(rest mass of photon is zero).Theory of relativity states that the faster you move the heavier you get.
In case of photon, if it moves with speed of light(which it does), it doesn’t have infinity mass.Why?
mv=m0/(1-v2/c2)^0.5 , where mvis mass of particle when the object is moving with ‘v’ velocity. m0is the rest mass.
For a photon moving with a speed of light:
v=c and mvtakes a form of 0/0. By applying a limits we get a definite mass.

As far as real mass of electron is concerned its the rest mass. That is one single electron which is at rest. v=0

Thanks & Regards
Aziz Alam
IIT Roorkee
askIITians.
shubham sharda
8 years ago
Mass of thoton depends upon the frame of refernce from which you are observing the photon. If your frame of reference is at rest with respect to the photon, then mass is zero.